suppose a firm has turned down an offer to pay 10,000 dollars to generate new sales of 100,000 dollars with a probability of .3. Would this firm be willing to take part in an expo by paying 5000 dollars, knowing that this will bring new sales of 50,000 dollars with a probability of .4?
Answer #1

profit in first case = 100,000 - 10,000 = 90000

expected profit in first case = 90000 * 0.3 = 27000

profit in second case = 50,000 - 5,000 = 45000

expected profit in second case = 45000 * 0.4 = 18000

the first case is profitable

press like